OpenStudy (anonymous):

Simplify the function f(t) = sinh(ln(t)). Then, find the derivative of your simplified form of f(t).

OpenStudy (shayaan_mustafa):

Hi ScottHann :) I think it is much easy.

OpenStudy (anonymous):

f(t)= t/2 - 1/2t

OpenStudy (kainui):

Sinh is the same as (e^x-e^-x)/2 and as we know from properties of logarithms e^lnx=x

OpenStudy (anonymous):

f'(t)= 1/2 + 1/2t^2

OpenStudy (shayaan_mustafa):

\[\Large\sinh=\frac{e^x-e^{-x}}{2}\] If \[\large x=ln(t)\] then\[\large e^x=t\] and \[\large e^{-x}=1/t\] So \[\Large sinh=\frac{t-1/t}{2}\] Now I think you can simplify it more.

OpenStudy (kainui):

Shayaan_Mustafa is right.

OpenStudy (shayaan_mustafa):

I have forgotten to use complete it should be look like this \[\Large sinh(lnt)=\frac{t-1/t}{2}\]

OpenStudy (shayaan_mustafa):

I forgot to use argument.

OpenStudy (anonymous):

Shayaan, thanks. I came up with (t^2-1)/2t, and Web Assign said it was wrong.

OpenStudy (shayaan_mustafa):

Well OK. Good luck.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Shayaan, thanks for your help. Your answer was correct.